# A Neurosymbolic Framework for Bias Correction in CNNs

### v2

@@ Abstract
Recent efforts in interpreting Convolutional Neural Networks (CNNs) focus on translating the activation of CNN filters into a stratified Answer Set Program (ASP) rule-sets. The CNN filters are known to capture high-level image concepts, thus the predicates in the rule-set are mapped to the concept that their corresponding filter represents. Hence, the rule-set effectively exemplifies the decision-making process of the CNN in terms of the concepts that it learns for any image classification task. These rule-sets help expose and understand the biases in CNNs, although correcting the biases effectively remains a challenge. We introduce a neurosymbolic framework called NeSyBiCor for bias correction in a trained CNN. Given symbolic concepts that the CNN is biased towards, expressed as ASP constraints, we convert the undesirable and desirable concepts to their corresponding vector representations. Then, the CNN is retrained using our novel semantic similarity loss that pushes the filters away from the representations of concepts that are undesirable while pushing them closer to the concepts that are desirable. The final ASP rule-set obtained after retraining, satisfies the constraints to a high degree, thus showing the revision in the knowledge of the CNN for the image classification task. We demonstrate that our NeSyBiCor framework successfully corrects the biases of CNNs trained with subsets of classes from the Places dataset while sacrificing minimal accuracy and improving interpretability, by greatly decreasing the size of the final bias-corrected rule-set w.r.t. the initial rule-set.
@@ 1 Introduction
Recent works have shown that it is possible to obtain the knowledge of a trained CNN in the form of a symbolic rule-set, more specifically as a stratified Answer Set Program (Padalkar et al., 2024a ; Padalkar et al., 2024b ). The authors proposed a framework called NeSyFOLD, wherein the activation of filters in the final convolutional layer of the CNN serves as the truth value of predicates in the generated rule-set, offering valuable insights into the concepts learnt by the model and their relation to the target class to be predicted. The filters in the CNN are $n\times n$ real-valued matrices. These matrices capture the representation of various concepts in the images. The predicates are labelled as the concept(s) that their corresponding filters learn to identify in the images. Fig. 1 illustrates the NeSyFOLD framework and the final rule-set that is generated for a train set containing images of the ‘bathroom’, ‘bedroom’ and ‘kitchen’ classes.
@@ 2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) are a sub-category of neural networks (NNs) well suited to pattern recognition in visual data first introduced by LeCun et al., (1989). Their distinctive feature is the presence of convolutional layers which employ learnable filters (also called kernels) to extract spatial hierarchies of features from input data. These filters are designed to detect specific features, such as edges or textures, by applying a mathematical operation called convolution. This operation involves sliding the filter over the image and computing the dot product of the filter values and the original pixel values in the image. The result of this process is a feature map, which is a new representation of the image emphasizing the detected features. Each filter in a layer can produce a distinct feature map, collectively forming a complex representation of the input that assists the network in learning to classify images or recognize patterns efficiently. It has been shown that when a CNN is trained with images, the filters in its convolutional layers learn specific patterns. Moreover, the last layer filters learn to recognize high-level features such as objects or object parts. This has been used to increase the explainability of CNNs via analysis of the types of emergent patterns and the relations between them (Zhang et al., (2018); Zhang et al., 2017a ). In this work, we build upon the line of research where CNN filters are used as symbolic atoms/predicates in a rule-set for image classification (Townsend et al., (2021); Padalkar et al., 2024a ; Padalkar et al., 2024b ).
@@ 2.2 FOLD-SE-M and NeSyFOLD
Padalkar et al., 2024a introduced a neurosymbolic framework called NeSyFOLD to extract a symbolic rule-set in the form of a stratified Answer Set Program from the last layer filters of a CNN. Fig. 1 illustrates the NeSyFOLD framework.
Binarization of filter outputs: For each image, the output feature maps of the last layer filters in the CNN are collected and stored. Next, the norms of the feature maps are computed, thus creating a table of norms of dimensions equal to no. of images $\times$ no. of filters. Finally, these norm values for each filter are treated as the filter’s activation value and a weighted sum of the mean and standard deviation of these values over all the rows (images) in the table determines the threshold of activation for each filter. This threshold is used for binarizing each value. This creates a binarized vector representation of the image with a dimension equal to the number of filters in the last layer. Such vectors are computed for all images in the training set and collected in a binarization table (Fig.1 top-right). Next, the FOLD-SE-M algorithm takes the binarization table as input and outputs a raw rule-set (Fig. 1 bottom-right) wherein each predicate represents a filter in the CNN’s last layer. The truth value of each predicate is determined by the binarized activation of the corresponding filter in the CNN when classifying a new image using the rule-set.
@@
Semantic labelling of predicates: Semantic segmentation masks of the images are used to map each filter to a set of concepts that it represents and thus the predicates in the rule-set can be given a semantic meaning. These semantic segmentation masks can be human annotated or those generated by using large foundation models such as SegGPT (Wang et al., (2023)) or RAM (Zhang et al., (2023)) in conjunction with the Segment Anything Model (SAM) (Kirillov et al., (2023)). Hence, the rule-set serves as a highly interpretable global explanation of the CNN’s decision making process. The predicted class of the image is represented in terms of logical rules based on concepts that the CNN filters learn individually. Fig. 2 illustrates the semantic labelling of a single predicate in the raw rule-set extracted by the FOLD-SE-M algorithm or in other words it maps the predicate in the raw rule-set with semantic concept(s) that its corresponding filter has learnt. The filters in the last convolutional layer are known to learn high-level concepts, such as objects or object parts (Zhou et. al @ICLR 2014). First, the feature maps generated by a filter in the last convolutional layer for the top-m images that activate it are collected. The top-m images are selected according to the norm values of the feature maps generated by the filter for all images in the train set. Next, the feature maps are resized and masked onto the top-m images (Fig. 2 top). Notice that by doing this, one can observe the concepts that the filter is looking at in each of selected images. For example, the filter considered in Fig. 2 is looking at beds in images. The masked images are then overlapped onto the semantic segmentation masks and an Intersection over Union (IoU) score is calculated for each concept that is visible after the overlap (Fig. 2, middle). Finally, the scores are aggregated over the top-m images and the label of the predicate that is associated with the filter under consideration is determined (Fig. 2, bottom).
@@
We refer the reader to the NeSyFOLD paper by Padalkar et al., 2024a for more details on the semantic labelling procedure and the NeSyFOLD framework itself.
@@ 4 Experiments
Result: In Fig. 4 we show the initial (RULE-SET 1 & RULE-SET 2) and final bias-corrected (RULE-SET 1* & RULE-SET 2*) rule-sets for $2$ of the $6$ subsets, namely des and defs. Due to the lack of space, we present the other rule-sets in the appendix.
@@
Finally, the most relevant observation is that the % Undesired value is always lesser for the bias-corrected rule-set as expected with an average reduction of $58\%$. This means that there is a reduction in the number of images being classified by following a path in the decision-making that involved an undesired concept. Since we used a higher $\lambda_{B}$ than $\lambda_{G}$ while calculating the semantic similarity loss, the CNN focused more on learning representations that are dissimilar to the undesired concepts. The % Desired value is higher in all cases except for babe and deh with an average increase of $35\%$. This means that the number of images being classified with at least one desirable concept in the decision path is increased. Note that in the case of babe and deh lower % Desired value means that the images were being classified by some other concepts that the human is indifferent towards. In practical scenarios, the primary concern is to ensure images are not categorized based on undesired concepts, a capability effectively demonstrated by the NeSyBiCor framework in our experiments.
@@ 6 Conclusion and Future Work
We proposed a neurosymbolic framework called NeSyBiCor for correcting the bias of a CNN. Our framework, in addition to correcting the bias of a CNN also allows the user to fine-tune the bias based on general concepts according to their specific needs or application. To the best of our knowledge this is the first method that does bias correction by using the learnt representations of the CNN filters in a targeted manner. We show through our experiments that the bias-corrected rule-set is highly effective at avoiding the classification of images based on undesired concepts. It is also more likely to classify the images based on the desired concepts. The main component of the NeSyBiCor framework is the semantic similarity loss. It serves as a measure of similarity between the representations learnt by the filters, to the representation of the undesired and desired concepts. Another benefit of the NeSyBiCor framework is that the bias-corrected rule-set is smaller in size, thus improving the interpretability.
@@
Finally, the work we presented here may be used to extend implementations of loss functions based on Differentiable Logics: Fischer et al., (2019); Ślusarz et al., (2023); Daggitt et al., (2024). So far, the idea of compiling a loss function from an arbitrary logical formula (advocated in the above papers) was based on the assumption that during training, checking for property satisfaction is as easy as checking class labels. And therefore, logical loss functions can be easily combined with the standard cross-entropy loss function in backpropagation training. For example, the famous property of neural network robustness involves a trivial robustness decision procedure at the training time (cf. Casadio et al., (2022)). This paper gives an example of a neuro-symbolic training scenario when a property of interest cannot be decided at the training time, and we had to resort to using CNN filters in order to replace a deterministic decision procedure. As the Differentiable Logic community continues to extend its range of applications (see e.g. Flinkow et al., (2024)), solutions such as the ones presented here will be needed in order to approximate the decision procedures. We intend to investigate this line of research.
Appendix A
In the main paper we could only show the initial and bias-corrected rule-sets for two of the six chosen subsets of the Places dataset. Hence, here we present the pre bias correction rule-sets and the bias-corrected rule-sets for the other four chosen subsets.
A.1 babek (‘bathroom’, ‘bedroom’, ‘kitchen’):
Initial:
target(X,‘kitchen’) :- cabinet1_wall1(X), not ab4(X), not ab5(X).
target(X,‘bedroom’) :- bed5(X), not ab6(X).
target(X,‘bathroom’) :- sink1_toilet1_wall4(X).
target(X,‘bathroom’) :- not cabinet4(X), wall10(X), not cabinet9_wall9(X).
target(X,‘bedroom’) :- bed9(X), not ab7(X), not ab9(X).
target(X,‘kitchen’) :- cabinet3(X).
target(X,‘bathroom’) :- sink2_wall5(X).
target(X,‘bathroom’) :- not bed1(X), not bed6(X), wall7_sink3(X).
target(X,‘bedroom’) :- bed4(X).
target(X,‘kitchen’) :- cabinet9_wall9(X).
target(X,‘bathroom’) :- not wall6(X), not bed7(X).
ab1(X) :- not bed6(X), cabinet3(X).
ab2(X) :- bed3(X), not ab1(X).
ab3(X) :- cabinet9_wall9(X), not ab2(X).
ab4(X) :- not cabinet5(X), not ab3(X).
ab5(X) :- not cabinet7(X), not cabinet6_wall8(X), sink1_toilet1_wall4(X).
ab6(X) :- wall7_sink3(X), sink1_toilet1_wall4(X).
ab7(X) :- not bed8(X), wall2(X).
ab8(X) :- not cabinet8(X), not cabinet2_wall3(X).
ab9(X) :- not bed2(X), not bed3(X), not ab8(X).
Bias-Corrected:
target(X,‘kitchen’) :- cabinet2_wall2(X), not ab1(X).
target(X,‘bedroom’) :- bed2(X).
target(X,‘bathroom’) :- sink1_toilet1(X).
target(X,‘bathroom’) :- not bed1(X), not ab2(X).
target(X,‘bedroom’) :- bed1(X).
ab1(X) :- not cabinet1(X), bed2(X).
ab2(X) :- not wall1(X), not wall3(X).
A.2 dedrh (‘desert road’, ‘driveway’, ‘highway’):
Initial:
target(X,‘highway’) :- not ground1(X), not ab4(X), not ab5(X).
target(X,‘driveway’) :- house3_building3(X), not ab7(X), not ab8(X).
target(X,‘desert_road’) :- not house2_sky2_trees2_building1_tree1_ground4_buildings1(X), not ab11(X), not ab12(X), not ab13(X).
target(X,‘highway’) :- road9(X), not ab14(X).
target(X,‘driveway’) :- house1(X).
target(X,‘highway’) :- road7(X).
ab1(X) :- not trees3(X), not road1(X).
ab2(X) :- not road15(X), trees3(X), not road13_car1(X), not sky1(X).
ab3(X) :- not house3_building3(X), not ab1(X), not ab2(X).
ab4(X) :- not road14(X), not ab3(X).
ab5(X) :- ground3(X), building2(X).
ab6(X) :- not road8(X),
    not road14(X), house2_sky2_trees2_building1_tree1_ground4_buildings1(X).
ab7(X) :- not house1(X), road5(X), not ab6(X).
ab8(X) :- not sky6_house4(X), sky5(X).
ab9(X) :- not trees1(X), sky4(X).
ab10(X) :- sky3(X), not ground2_road4(X), not house5_building4(X).
ab11(X) :- not mountain1(X), not road3(X), not ab9(X), not ab10(X).
ab12(X) :- road10(X), not road11(X).
ab13(X) :- road1(X), trees3(X), not road12(X).
ab14(X) :- not road6(X), road2(X).
Bias-Corrected:
target(X,‘highway’) :- not ground1_road2(X), not house1(X), not sky2(X), not ab1(X).
target(X,‘driveway’) :- not sky2(X), not ab2(X), not ab3(X), not ab4(X).
target(X,‘desert_road’) :- sky2(X).
target(X,‘desert_road’) :- ground1_road2(X), not trees1(X), not ab5(X).
ab1(X) :- not road3(X), building1_house3_road5_car1(X).
ab2(X) :- not house1(X), not building1_house3_road5_car1(X), not house2_trees2(X).
ab3(X) :- road3(X), road1(X).
ab4(X) :- not house1(X), road4(X), not tree1_sky1(X).
ab5(X) :- not mountain1(X), grass1_trees3(X).
A.3 babe (‘bathroom’, ‘bedroom’):
Initial:
target(X,‘bedroom’) :- not mirror1(X,’1’), not bed4(X,’0’).
target(X,‘bathroom’) :- not countertop1(X,’0’).
target(X,‘bedroom’) :- not bed7(X,’0’), not ab1(X,’True’).
target(X,‘bathroom’) :- bed3(X,’0’).
target(X,‘bedroom’) :- bed3(X,’0’).
target(X,‘bedroom’) :- not bathtub1(X,’1’), bed6(X,’1’).
target(X,‘bedroom’) :- bed2(X,’1’).
target(X,‘bathroom’) :- mirror1(X,’1’).
ab1(X,’True’) :- not bed1(X,’1’), not bathtub1(X,’0’), not bed5(X,’1’).
Bias-Corrected:
target(X,‘bedroom’) :- bed1(X).
target(X,‘bathroom’) :- not bed3(X), not ab1(X).
target(X,‘bedroom’) :- not sink1(X).
ab1(X) :- not sink1(X), bed2(X).
A.4 deh (‘desert road’, ‘highway’):
Initial:
target(X,‘highway’) :- sky2_trees1(X), not ab1(X), not ab2(X).
target(X,‘desert_road’) :- ground1(X), not ab3(X).
target(X,‘highway’) :- road3(X), not ab4(X).
target(X,‘desert_road’) :- sky1_road2(X), not ab5(X).
target(X,‘highway’) :- sky3(X).
target(X,‘desert_road’) :- road4(X).
target(X,‘highway’) :- road1(X).
target(X,‘desert_road’) :- not road7(X).
ab1(X) :- not road3(X), not trees2(X), sky4(X).
ab2(X) :- sky1_road2(X), ground1(X), not road5(X).
ab3(X) :- road9_trees3(X), road6(X), road13(X).
ab4(X) :- sky1_road2(X), not road10(X).
ab5(X) :- road11(X), road12(X), not road8(X).
Bias-Corrected:
target(X,‘highway’) :- trees2(X).
target(X,‘desert_road’) :- ground1(X).
target(X,‘desert_road’) :- not road4(X).
target(X,‘highway’) :- not road1(X).
target(X,‘highway’) :- not road2(X).

### v3

@@ Abstract
Recent efforts in interpreting Convolutional Neural Networks (CNNs) focus on translating the activation of CNN filters into a stratified Answer Set Program (ASP) rule-sets. The CNN filters are known to capture high-level image concepts, thus the predicates in the rule-set are mapped to the concept that their corresponding filter represents. Hence, the rule-set exemplifies the decision-making process of the CNN w.r.t the concepts that it learns for any image classification task. These rule-sets help understand the biases in CNNs, although correcting the biases remains a challenge. We introduce a neurosymbolic framework called NeSyBiCor for bias correction in a trained CNN. Given symbolic concepts, as ASP constraints, that the CNN is biased towards, we convert the concepts to their corresponding vector representations. Then, the CNN is retrained using our novel semantic similarity loss that pushes the filters away from (or towards) learning the desired/undesired concepts. The final ASP rule-set obtained after retraining, satisfies the constraints to a high degree, thus showing the revision in the knowledge of the CNN. We demonstrate that our NeSyBiCor framework successfully corrects the biases of CNNs trained with subsets of classes from the Places dataset while sacrificing minimal accuracy and improving interpretability.
@@ 1 Introduction
Recent works have shown that it is possible to obtain the knowledge of a trained CNN in the form of a symbolic rule-set, more specifically as a stratified Answer Set Program (Padalkar et al., 2024b ; Padalkar et al., 2024c ). The authors proposed a framework called NeSyFOLD, wherein the activation of filters in the final convolutional layer of the CNN serves as the truth value of predicates in the generated rule-set, offering valuable insights into the concepts learnt by the model and their relation to the target class to be predicted. The filters in the CNN are $n\times n$ real-valued matrices. These matrices capture the representation of various concepts in the images. The predicates are labelled as the concept(s) that their corresponding filters learn to identify in the images. Fig. 1 illustrates the NeSyFOLD framework and the final rule-set that is generated for a train set containing images of the ‘bathroom’, ‘bedroom’ and ‘kitchen’ classes.
@@ 2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) are a sub-category of neural networks (NNs) well suited to pattern recognition in visual data first introduced by LeCun et al., (1989). Their distinctive feature is the presence of convolutional layers which employ learnable filters (also called kernels) to extract spatial hierarchies of features from input data. These filters are designed to detect specific features, such as edges or textures, by applying a mathematical operation called convolution. This operation involves sliding the filter over the image and computing the dot product of the filter values and the original pixel values in the image. The result of this process is a feature map, which is a new representation of the image emphasizing the detected features. Each filter in a layer can produce a distinct feature map, collectively forming a complex representation of the input that assists the network in learning to classify images or recognize patterns efficiently. It has been shown that when a CNN is trained with images, the filters in its convolutional layers learn specific patterns. Moreover, the last layer filters learn to recognize high-level features such as objects or object parts. This has been used to increase the explainability of CNNs via analysis of the types of emergent patterns and the relations between them (Zhang et al., (2018); Zhang et al., 2017a ). In this work, we build upon the line of research where CNN filters are used as symbolic atoms/predicates in a rule-set for image classification (Townsend et al., (2021); Padalkar et al., 2024b ; Padalkar et al., 2024c ).
@@ 2.2 FOLD-SE-M and NeSyFOLD
Padalkar et al., 2024b introduced a neurosymbolic framework called NeSyFOLD to extract a symbolic rule-set in the form of a stratified Answer Set Program from the last layer filters of a CNN. Fig. 1 illustrates the NeSyFOLD framework.
Binarization of filter outputs: For each image, the output feature maps of the last layer filters in the CNN are collected and stored. Next, the norms of the feature maps are computed. A norm of a matrix here is simply adding the squared values of each element in the matrix and taking a square root of the output. Eq. 2 depicts this operation for the feature map produced by the $k^{th}$ filter for the $i^{th}$ image. This creates a table of norms of dimensions equal to no. of images $\times$ no. of filters. Finally, these norm values for each filter are treated as the filter’s activation value and a weighted sum of the mean and standard deviation (Eq. 3) of these values over all the rows (images) in the table determines the threshold of activation for each filter. $\alpha$ and $\gamma$ are hyperparameters in Eq. 3 and $n$ is the number of images in the train-set. $\theta_{k}$ is the threshold for the $k^{th}$ filter. This threshold is used to binarize each value. This creates a binarized vector representation of the image with a dimension equal to the number of filters in the last layer. Such vectors are computed for all images in the training set and collected in a binarization table (Fig.1 top-right). Next, the FOLD-SE-M algorithm takes the binarization table as input and outputs a raw rule-set (Fig. 1 bottom-right) wherein each predicate represents a filter in the CNN’s last layer. The truth value of each predicate is determined by the binarized activation of the corresponding filter in the CNN when classifying a new image using the rule-set.
$a_{i,k}=$ $~{}||A_{i,k}||_{2}$ (2)
$\theta_{k}=$ $~{}\alpha\cdot\overline{a_{k}}+\gamma\sqrt{\frac{1}{n}\sum(a_{i,k}-\overline{a_{k}})^{2}}$ (3)
@@
Semantic labelling of predicates: Semantic segmentation masks of the images are used to map each filter to a set of concepts that it represents and thus the predicates in the rule-set can be given a semantic meaning. These semantic segmentation masks can be human annotated or those generated by using large foundation models such as SegGPT (Wang et al., (2023)) or RAM (Zhang et al., (2023)) in conjunction with the Segment Anything Model (SAM) (Kirillov et al., (2023)). Hence, the rule-set serves as a highly interpretable global explanation of the CNN’s decision making process. The predicted class of the image is represented in terms of logical rules based on concepts that the CNN filters learn individually. Fig. 2 illustrates the semantic labelling of a single predicate in the raw rule-set extracted by the FOLD-SE-M algorithm or, in other words, it maps the predicate in the raw rule-set with semantic concept(s) that its corresponding filter has learnt. The filters in the last convolutional layer are known to learn high-level concepts, such as objects or object parts Zhou et al., (2015). First, the feature maps generated by a filter in the last convolutional layer for the top-m images that activate it are collected. The top-m images are selected according to the norm values of the feature maps generated by the filter for all images in the train set. Next, the feature maps are resized and masked onto the top-m images (Fig. 2 top). Notice that by doing this, one can observe the concepts that the filter is looking at in each of selected images. For example, the filter considered in Fig. 2 is looking at beds in images. The masked images are then overlapped onto the semantic segmentation masks and an Intersection over Union (IoU) score is calculated for each concept that is visible after the overlap (Fig. 2, middle). Finally, the scores are aggregated over the top-m images and the label of the predicate that is associated with the filter under consideration is determined (Fig. 2, bottom).
@@
We refer the reader to the NeSyFOLD paper by Padalkar et al., 2024b for more details on the semantic labelling procedure and the NeSyFOLD framework itself.
@@ 4 Experiments
Result: In Fig. 4 we show the initial (RULE-SET 1 & RULE-SET 2) and final bias-corrected (RULE-SET 1* & RULE-SET 2*) rule-sets for $2$ of the $6$ subsets, namely des and defs. Due to the lack of space, we present the other rule-sets in the supplementary material (Padalkar et al., 2024a ).
@@
Finally, the most relevant observation is that the % Undesired value is always lesser for the bias-corrected rule-set as expected with an average reduction of $58\%$. This means that there is a reduction in the number of images being classified by following a path in the decision-making that involved an undesired concept. Since we used a higher $\lambda_{B}$ than $\lambda_{G}$ while calculating the semantic similarity loss, the CNN focused more on learning representations that are dissimilar to the undesired concepts. The % Desired value is higher in all cases except for babe and deh with an average increase of $35\%$. This means that the number of images being classified with at least one desirable concept in the decision path is increased. Note that in the case of babe and deh, a lower % Desired value means that the images were being classified by some other concepts that the human is indifferent towards. In practical scenarios, the primary concern is to ensure images are not categorized based on undesired concepts, a capability effectively demonstrated by the NeSyBiCor framework in our experiments.
@@ 6 Conclusion and Future Work
We proposed a neurosymbolic framework called NeSyBiCor for correcting the bias of a CNN. Our framework, in addition to correcting the bias of a CNN also allows the user to fine-tune the bias based on general concepts according to their specific needs or application. To the best of our knowledge, this is the first method that does bias correction by using the learnt representations of the CNN filters in a targeted manner. We show through our experiments that the bias-corrected rule-set is highly effective at avoiding the classification of images based on undesired concepts. It is also more likely to classify the images based on the desired concepts. The main component of the NeSyBiCor framework is the semantic similarity loss. It serves as a measure of similarity between the representations learnt by the filters, to the representation of the undesired and desired concepts. Another benefit of the NeSyBiCor framework is that the bias-corrected rule-set is smaller in size, thus improving the interpretability.
@@
Finally, the work we presented here may be used to extend implementations of loss functions based on Differentiable Logics: Fischer et al., (2019); Ślusarz et al., (2023); Daggitt et al., (2024). So far, the idea of compiling a loss function from an arbitrary logical formula (advocated in the above papers) was based on the assumption that during training, checking for property satisfaction is as easy as checking class labels. Therefore, logical loss functions can be easily combined with the standard cross-entropy loss function in backpropagation training. For example, the famous property of neural network robustness involves a trivial robustness decision procedure at the training time (cf. Casadio et al., (2022)). This paper gives an example of a neuro-symbolic training scenario when a property of interest cannot be decided at the training time, and we had to resort to using CNN filters in order to replace a deterministic decision procedure. As the Differentiable Logic community continues to extend its range of applications (see e.g. Flinkow et al., (2024)), solutions such as the ones presented here will be needed in order to approximate the decision procedures. We intend to investigate this line of research.